OpenStudy (anonymous):

what is 5*5111*548762*/154454659=?

OpenStudy (anonymous):

use a calculator

OpenStudy (anonymous):

(5*5111*548762)/154454659 = 14023612910/154454659 = approx. 90.79437

OpenStudy (anonymous):

Use Operations order .

OpenStudy (anonymous):

5*5111*548762*/154454659 paste to google

OpenStudy (ja1):

First we must use order of operations

OpenStudy (ja1):

what is 5*5111

OpenStudy (ja1):

Search "Calculator" in google to help you

OpenStudy (ja1):

5*5111 = 25555 2555*548762 = 1402086910 1402086910/154454659 = 9.07766019541 Answer = 9.07766019541

OpenStudy (ja1):

@lovekins13 there's your answer :)

OpenStudy (ja1):

I hope I did that right xD

OpenStudy (goformit100):

for doing this question two way you can solve this question. first Mentally Second Using calculator which you prefer ?

OpenStudy (goformit100):

if you do it using calculator it seems easy. But by doing mentally you will have to proceed as follows.

OpenStudy (goformit100):

(5*5111*548762) = 14023612910 = A(say)

OpenStudy (goformit100):

154454659 = B(Say)

OpenStudy (goformit100):